OpenStudy (anonymous):

the relationship between fahrenheit and celsius temperature can be approximated by and equation of the form y=mx + b, where x is the temperature in degrees fahrenheit and y is the temperature in degrees celsius . find the equation if two ordered pairs that satify it are (50,10) and (86,30).

OpenStudy (mertsj):

Slope = 20/36=5/9 y-10=5/9(x-50) y-10=5/9x-250/9 y=5/9x-160/9

OpenStudy (aravindg):

ys